Can someone please help me with this problem?
The price of a doll increased from $600 to $800. What was the percent of the increase? Round to the nearest tenth percent.
Thanks in advance for your help!

Answer by stanbon(75887)

The price of a doll increased from $600 to $800. What was the percent of the increase? Round to the nearest tenth percent.
---------------
Quantity increase: 800-600= $200
----------
Percent increase: 200/600 = 1/3 = 33.3%
==========
Cheers,
Stan H.